OpenStudy (anonymous):

Determine the critical values: 2x^2-64sqrt(x)

OpenStudy (kainui):

Critical values occur when the slope is zero. Batman can take it away if he so desires! =P

OpenStudy (anonymous):

ok i know that but the square root is throwing me

OpenStudy (kainui):

Show me your best attempt. Also, you can rewrite a square root as:\[\sqrt{x}=x^{1/2}\] Which makes sense with the rest of exponent rules. =)

OpenStudy (anonymous):

i got that but then what?

OpenStudy (kainui):

Show me what you got.

OpenStudy (anonymous):

Do you know the definition of a critical value?

OpenStudy (anonymous):

yes the derivative = 0 I got the derivative: 4x-32sqrtx/x

OpenStudy (anonymous):

i just dont know how to slove for 0

OpenStudy (anonymous):

ok, here is the definition of a critical value. c is call a critical value of a function f if f'(c) = 0 or f'(c) does not exist. So there are two things you need to find: where f has slope = 0, or where f' does not exist. First thing is you need to take the derivative, which you got it incorrect f'(x) = 4x - 32x^(-1/2)

OpenStudy (anonymous):

oh yeah i just rationalized the denominator

OpenStudy (anonymous):

:o oh, didn't realize that. Ok, so x = 0 is certain one of the critical value since f'(0) does not exist. the other one is f'(x) = 0 4x-32sqrt(x)/x = 0 then solve for x.

OpenStudy (anonymous):

https://www.wolframalpha.com/input/?i=4x-32sqrt%28x%29%2Fx+%3D+0 so x = 0 or 4 <=== ans

OpenStudy (anonymous):

so how did ya get x=0?

OpenStudy (anonymous):

does f'(0) exist?

OpenStudy (anonymous):

in other words, can you do f'(0) = 4x - 32sqrt(0)/0 ?

OpenStudy (anonymous):

C is also in the domain of f(x) not sure if you mentioned that, but just pointing it out.

OpenStudy (anonymous):

but how does the derivative=0 come out to be 0? I got the 4, but not the 0

OpenStudy (anonymous):

Set your derivative = 0

OpenStudy (anonymous):

You have to check where f'(x) = 0 and where f'(x) = undefined.

OpenStudy (anonymous):

x = 4 makes f' = 0. But the fact that x = 0 is also a critical value is because f'(0) does not exist. and 0 is also in the domain of f as iambatman mentioned

OpenStudy (anonymous):

So if you can show your work, it'll be easier to guide you.

OpenStudy (anonymous):

ok so4x-32x^.5=0 4x-32/sqrtx=0 4x=32/sqrtx 32=4xsqrtx 8=x^3/2 x=4 but i dont see how it can be 0 from this anyways :)

OpenStudy (anonymous):

again, if you look at the definition, 0 didn't come out from solving for x of the equation f'(x) = 0. It comes from the fact that 0 is in the domain of f, and f'(0) does not exist.

OpenStudy (anonymous):

ok so did i solve for 0 right?

OpenStudy (anonymous):

you mean if you found x = 0 the right way?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

you only found x = 4.

OpenStudy (anonymous):

so i guess my question is,, is there a way to find the x=o like i found the x=4? sorry i confused you

OpenStudy (anonymous):

You get 0 from seeing where f'(x) is undefined.

OpenStudy (anonymous):

yes. Most of the times, it comes from the denominator because the's where division by 0 occurs. recall that f'(x) = 4x - 32sqrt(x)/x. Whatever the denominator is, you set it equals to 0 and solve. since x is is the only term that is in the denominator, setting it equal to 0, i.e x = 0, which gives you x = 0 when you solve

OpenStudy (anonymous):

oh ok i see it halleluja! haha thanks!!! both of you!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

These are the main things you should remember when looking for critical points. f'(x) = 0 f'(x) = undefined c is in the domain of f(x) If I'm missing something please let me know @sourwing XD

OpenStudy (anonymous):

@iambatman that's correct. those are the only 3 criteria to find the critical values.

OpenStudy (anonymous):

Thought so, just checking :p